OpenStudy (anonymous):

Please... i need the answer and how you get the answer... The perimeter of a rectangle is twice the sum of its length and width. The perimeter is 28 meters and its length is 2 meters more than twice its width. What is the length. Thank you in advance

OpenStudy (anonymous):

P = 2l + 2w l = 2 + w P = 2(2 + w) + 2w 28 = 4 + 2w + 2w 28 = 4 + 4w 24 = 4w w = 6 l = 2 + 6 = 8

OpenStudy (anonymous):

mkuehn... i can't grasp your concept/formula... i'm faced with a similar question the last answer was incorrect and I don't know how to figure it out. what is the formula in the simplest terms possible. (pretend I'm 2) the perimeter of a rectangle is twice the sum of its length and width. The perimeter is 34 meters and its length is 2 meters more than twice its width. What is its length? (and how)

OpenStudy (anonymous):

Okay so the perimeter of a rectangle is twice the sum of its length and width. This means: P = 2(l + w) = 2l + 2w We need to write a relationship between the length and the width. We are told that the length is 2 meters more than twice its width. l = 2 + 2w Now substitute this back into the perimeter formula and solve for the width. P = 2(2 + 2w) + 2w 34 = 4 + 4w + 2w 34 = 4 + 6w 30 = 6w w = 5 Now we know the width so we can use our relationship with the length to figure that out. l = 2 + 2w = 2 + 2(5) = 12 The rectangle has length 12 and width 5. Its perimeter is then 12 + 12 + 5 + 5 = 34. 2 more than twice its width is 2 + 2(5) = 12.

OpenStudy (anonymous):

thank you. I do believe I will just be an idiot in this subject or eternity. I just don't get how you arrive at the answer... I can't put it together

OpenStudy (anonymous):

It's all about translating the English into math. When you see something like "length is 2 meters more than twice its width" you have to translate that as being length = 2 + 2 x width Since there are two unknowns (length and width), you need to rewrite the perimeter formula in terms of one of these variables and solve for it. Then, using the relationship between the two, solve for the other variable.

OpenStudy (anonymous):

i'm laughing.... this made it so much easier to understand. (not really, I'm just going to have live with being and idiot... again!!) thank you, really. I appreciate your effort. I've got a brick for a brain